OpenStudy (anonymous):

-x+2y=8 the x-value when y=3 is x=

OpenStudy (anonymous):

you can substitute the values in

OpenStudy (anonymous):

@mth3v4 im lost

OpenStudy (anonymous):

what does y =

OpenStudy (anonymous):

3@mth3v4

OpenStudy (anonymous):

@mth3v4

OpenStudy (anonymous):

ok can you perform the indicated operations

OpenStudy (anonymous):

im still wondering how it =3, i was thinking -10+2=8 but i know that aint right now

OpenStudy (anonymous):

sry not -10 but idk

OpenStudy (anonymous):

it is a given it said y= 3 2( 3)

OpenStudy (texaschic101):

-x + 2(3) = 8 can you solve this ??

OpenStudy (anonymous):

@texaschic101 is it -2+2(3)=8

OpenStudy (texaschic101):

where are you getting the -2 from ?

OpenStudy (anonymous):

idk

OpenStudy (texaschic101):

-x + 2(3) = 8 -x + 6 = 8 -x = 8 - 6 -x = 2 x = -2 check... -(-2) + 6 = 8 2 + 6 = 8 8 = 8 (correct) My bad.....-2 = x.......you are correct

OpenStudy (anonymous):

@mth3v4 is this correct

OpenStudy (texaschic101):

lets check it......x - 2, y = 3 -x + 2y = 8 -(-2) + 2(3) = 8 2 + 6 = 8 8 = 8 (correct) Yes it is correct

OpenStudy (texaschic101):

typo...x = - 2

OpenStudy (anonymous):

@texaschic101 i have 1 more ? its another problem though, im about to post it

OpenStudy (anonymous):

\[-1/4t \ge1\] the solution set is??

OpenStudy (anonymous):

btw could you close this post and post new questions on a diff post :P

OpenStudy (anonymous):

it reduces clutter

OpenStudy (texaschic101):

@mth3v4 is correct....1 question per post is better

OpenStudy (anonymous):

yes